A deck (52) of cards (clubs, diamonds, hearts and spades) is shuffled once. Estimate the probability of drawing a king and queen. Estimate the probability of drawing a 4 (four). Estimate the probability of drawing a card that is not a heart.

1
Expert's answer
2022-05-05T10:32:39-0400

It is not clear how many cards are drawn at first time. Therefore, we consider two possible cases:

  1. One card is drawn. The deck contains "8" kings and queens. Thus, the probability is: "p=\frac{8}{52}=\frac{2}{13}"
  2. Two cards are drawn. There are "C_{52}^2=\frac{52\cdot51}{2}=1326" different combinations. There are "C_8^2=\frac{8\cdot7}2=28" combinations that contain a king and a queen. The probability to get a king and a queen is: "\frac{28}{1326}\approx0.02". (it is rounded to 2 decimal places)

For all other cases we assume that there is only one card. The probability of getting "4" is: "\frac{4}{52}=\frac{1}{13}". There are "52-13=39" cards that are not heart. The probability to get not a heart is: "\frac{39}{52}=\frac{3}{4}".

Answer: The probability to get a king and a queen is: "\frac{2}{13}" for one card and "0.02" for two cards (it is rounded to two decimal places). The probability to get "4" is: "\frac{1}{13}". The probability to get not a heart is: "\frac{3}{4}"